OpenStudy (anonymous):

A car travels at an average speed of 52 miles per hour. How many miles does it travel in 4 hours and 45 minutes?

OpenStudy (lgbasallote):

4 hours and 45 minutes = \[4 \frac{45}{60} \; \text{hours}\] \[\implies 4 \frac{3}{4} \; \text{hours}\] now to find distance we use \[\text{distance} = \text{speed} \times \text{time}\] \[\implies 52 \times 4 \frac 34\]

OpenStudy (lgbasallote):

i guess you can turn \[ 4 \frac 34 \;\;\;\; \text{into improper first}\] \[4 \frac 34 = \frac{19}{4}\] \[52 \times \frac{19}{4}\] solve this now and you'll get distance

OpenStudy (anonymous):

Would it be 247 ?

OpenStudy (lgbasallote):

i dont have a calculator so idk heh

OpenStudy (anonymous):

247 miles

OpenStudy (anonymous):

oh. haha. well thanks for the help!